# Introducing Routing Functions to Vision-Language Parameter-Efficient Fine-Tuning with Low-Rank Bottlenecks

### v1

@@ 4.3 Results
$x_{t}\circ x^{\prime}_{v}$ $7.8\mathbf{26.1_{\color[rgb]{0.1914,0.3359375,0.9414}{+7.8}}}$ $6.3\mathbf{23.7_{\color[rgb]{0.1914,0.3359375,0.9414}{+6.3}}}$ $11.8\mathbf{48.6_{\color[rgb]{0.1914,0.3359375,0.9414}{+11.8}}}$ $32.9\mathbf{88.7_{\color[rgb]{0.1914,0.3359375,0.9414}{+32.9}}}$ $5.5\mathbf{17.3_{\color[rgb]{0.1914,0.3359375,0.9414}{+5.5}}}$ $12.86\mathbf{40.88_{\color[rgb]{0.1914,0.3359375,0.9414}{+12.86}}}$ $4.4\underline{25.4_{\color[rgb]{0.1914,0.3359375,0.9414}{+4.4}}}$ $2.8\mathbf{23.7_{\color[rgb]{0.1914,0.3359375,0.9414}{+2.8}}}$ $4.8\mathbf{48.2_{\color[rgb]{0.1914,0.3359375,0.9414}{+4.8}}}$ $17.6\mathbf{87.9_{\color[rgb]{0.1914,0.3359375,0.9414}{+17.6}}}$ $3.0\mathbf{17.7_{\color[rgb]{0.1914,0.3359375,0.9414}{+3.0}}}$ $6.52\mathbf{40.58_{\color[rgb]{0.1914,0.3359375,0.9414}{+6.52}}}$
$x_{t}+x^{\prime}_{v}$ $3.922.2_{\color[rgb]{0.1914,0.3359375,0.9414}{+3.9}}$ $2.920.3_{\color[rgb]{0.1914,0.3359375,0.9414}{+2.9}}$ $4.241.0_{\color[rgb]{0.1914,0.3359375,0.9414}{+4.2}}$ $17.773.5_{\color[rgb]{0.1914,0.3359375,0.9414}{+17.7}}$ $3.014.8_{\color[rgb]{0.1914,0.3359375,0.9414}{+3.0}}$ $6.3434.36_{\color[rgb]{0.1914,0.3359375,0.9414}{+6.34}}$ $4.4\underline{25.4_{\color[rgb]{0.1914,0.3359375,0.9414}{+4.4}}}$ $2.2\underline{23.1_{\color[rgb]{0.1914,0.3359375,0.9414}{+2.2}}}$ $4.7\underline{47.1_{\color[rgb]{0.1914,0.3359375,0.9414}{+4.7}}}$ $16.8\underline{87.1_{\color[rgb]{0.1914,0.3359375,0.9414}{+16.8}}}$ $2.317.0_{\color[rgb]{0.1914,0.3359375,0.9414}{+2.3}}$ $5.88\underline{39.94_{\color[rgb]{0.1914,0.3359375,0.9414}{+5.88}}}$
$x_{t}(x_{v})^{\mathsf{T}}x_{v}$ $6.5\underline{24.8_{\color[rgb]{0.1914,0.3359375,0.9414}{+6.5}}}$ $5.2\underline{22.6_{\color[rgb]{0.1914,0.3359375,0.9414}{+5.2}}}$ $8.6\underline{45.4_{\color[rgb]{0.1914,0.3359375,0.9414}{+8.6}}}$ $\underline{84.9_{+29.1}}$ $5.0\underline{16.8_{\color[rgb]{0.1914,0.3359375,0.9414}{+5.0}}}$ $10.88\underline{38.90_{\color[rgb]{0.1914,0.3359375,0.9414}{+10.88}}}$ $3.924.9_{\color[rgb]{0.1914,0.3359375,0.9414}{+3.9}}$ $1.822.7_{\color[rgb]{0.1914,0.3359375,0.9414}{+1.8}}$ $2.045.4_{\color[rgb]{0.1914,0.3359375,0.9414}{+2.0}}$ $13.583.8_{\color[rgb]{0.1914,0.3359375,0.9414}{+13.5}}$ $2.216.9_{\color[rgb]{0.1914,0.3359375,0.9414}{+2.2}}$ $4.6838.74_{\color[rgb]{0.1914,0.3359375,0.9414}{+4.68}}$
$x_{t}x^{\prime\prime}_{v}$ $5.623.9_{\color[rgb]{0.1914,0.3359375,0.9414}{+5.6}}$ $4.521.9_{\color[rgb]{0.1914,0.3359375,0.9414}{+4.5}}$ $7.143.9_{\color[rgb]{0.1914,0.3359375,0.9414}{+7.1}}$ $14.780.5_{\color[rgb]{0.1914,0.3359375,0.9414}{+14.7}}$ $4.416.2_{\color[rgb]{0.1914,0.3359375,0.9414}{+4.4}}$ $9.2637.28_{\color[rgb]{0.1914,0.3359375,0.9414}{+9.26}}$ $4.5\mathbf{25.5_{\color[rgb]{0.1914,0.3359375,0.9414}{+4.5}}}$ $2.022.9_{\color[rgb]{0.1914,0.3359375,0.9414}{+2.0}}$ $2.646.0_{\color[rgb]{0.1914,0.3359375,0.9414}{+2.6}}$ $14.985.2_{\color[rgb]{0.1914,0.3359375,0.9414}{+14.9}}$ $2.5\underline{17.2_{\color[rgb]{0.1914,0.3359375,0.9414}{+2.5}}}$ $5.3039.36_{\color[rgb]{0.1914,0.3359375,0.9414}{+5.30}}$
@@
$x_{t}\circ x^{\prime}_{v}$ $8.7\underline{24.6_{\color[rgb]{0.1914,0.3359375,0.9414}{+8.7}}}$ $4.6\underline{23.1_{\color[rgb]{0.1914,0.3359375,0.9414}{+4.6}}}$ $9.446.4_{\color[rgb]{0.1914,0.3359375,0.9414}{+9.4}}$ $22.984.5_{\color[rgb]{0.1914,0.3359375,0.9414}{+22.9}}$ $3.1\underline{17.2_{\color[rgb]{0.1914,0.3359375,0.9414}{+3.1}}}$ $9.7439.16_{\color[rgb]{0.1914,0.3359375,0.9414}{+9.74}}$ $7.523.6_{\color[rgb]{0.1914,0.3359375,0.9414}{+7.5}}$ $4.2\underline{22.6_{\color[rgb]{0.1914,0.3359375,0.9414}{+4.2}}}$ $8.6\underline{44.8_{\color[rgb]{0.1914,0.3359375,0.9414}{+8.6}}}$ $21.3\underline{82.8_{\color[rgb]{0.1914,0.3359375,0.9414}{+21.3}}}$ $3.2\underline{17.1_{\color[rgb]{0.1914,0.3359375,0.9414}{+3.2}}}$ $8.96\underline{38.18_{\color[rgb]{0.1914,0.3359375,0.9414}{+8.96}}}$
$x_{t}+x^{\prime}_{v}$ $5.121.0_{\color[rgb]{0.1914,0.3359375,0.9414}{+5.1}}$ $3.021.5_{\color[rgb]{0.1914,0.3359375,0.9414}{+3.0}}$ $5.442.4_{\color[rgb]{0.1914,0.3359375,0.9414}{+5.4}}$ $14.475.0_{\color[rgb]{0.1914,0.3359375,0.9414}{+14.4}}$ $1.916.0_{\color[rgb]{0.1914,0.3359375,0.9414}{+1.9}}$ $5.7635.18_{\color[rgb]{0.1914,0.3359375,0.9414}{+5.76}}$ $3.219.3_{\color[rgb]{0.1914,0.3359375,0.9414}{+3.2}}$ $2.420.8_{\color[rgb]{0.1914,0.3359375,0.9414}{+2.4}}$ $4.740.9_{\color[rgb]{0.1914,0.3359375,0.9414}{+4.7}}$ $13.875.2_{\color[rgb]{0.1914,0.3359375,0.9414}{+13.8}}$ $2.316.2_{\color[rgb]{0.1914,0.3359375,0.9414}{+2.3}}$ $5.2634.48_{\color[rgb]{0.1914,0.3359375,0.9414}{+5.26}}$
$x_{t}(x_{v})^{\mathsf{T}}x_{v}$ $\underline{26.1_{+10.2}}$ $4.7\mathbf{23.2_{\color[rgb]{0.1914,0.3359375,0.9414}{+4.7}}}$ $9.9\mathbf{46.9_{\color[rgb]{0.1914,0.3359375,0.9414}{+9.9}}}$ $23.8\underline{85.4_{\color[rgb]{0.1914,0.3359375,0.9414}{+23.8}}}$ $3.2\mathbf{17.3_{\color[rgb]{0.1914,0.3359375,0.9414}{+3.2}}}$ $10.36\underline{39.78_{\color[rgb]{0.1914,0.3359375,0.9414}{+10.36}}}$ $8.0\underline{24.1_{\color[rgb]{0.1914,0.3359375,0.9414}{+8.0}}}$ $4.022.4_{\color[rgb]{0.1914,0.3359375,0.9414}{+4.0}}$ $8.044.2_{\color[rgb]{0.1914,0.3359375,0.9414}{+8.0}}$ $21.282.7_{\color[rgb]{0.1914,0.3359375,0.9414}{+21.2}}$ $3.016.9_{\color[rgb]{0.1914,0.3359375,0.9414}{+3.0}}$ $8.8438.06_{\color[rgb]{0.1914,0.3359375,0.9414}{+8.84}}$
$x_{t}x^{\prime\prime}_{v}$ $10.7\mathbf{26.6_{\color[rgb]{0.1914,0.3359375,0.9414}{+10.7}}}$ $4.523.0_{\color[rgb]{0.1914,0.3359375,0.9414}{+4.5}}$ $9.8\underline{46.8_{\color[rgb]{0.1914,0.3359375,0.9414}{+9.8}}}$ $24.2\mathbf{85.8_{\color[rgb]{0.1914,0.3359375,0.9414}{+24.2}}}$ $3.1\underline{17.2_{\color[rgb]{0.1914,0.3359375,0.9414}{+3.1}}}$ $10.46\mathbf{39.88_{\color[rgb]{0.1914,0.3359375,0.9414}{+10.46}}}$ $9.9\mathbf{26.0_{\color[rgb]{0.1914,0.3359375,0.9414}{+9.9}}}$ $5.1\mathbf{23.5_{\color[rgb]{0.1914,0.3359375,0.9414}{+5.1}}}$ $10.6\mathbf{46.8_{\color[rgb]{0.1914,0.3359375,0.9414}{+10.6}}}$ $24.8\mathbf{86.3_{\color[rgb]{0.1914,0.3359375,0.9414}{+24.8}}}$ $3.9\mathbf{17.8_{\color[rgb]{0.1914,0.3359375,0.9414}{+3.9}}}$ $10.86\mathbf{40.08_{\color[rgb]{0.1914,0.3359375,0.9414}{+10.86}}}$
@@
LoRA $x_{t}(x_{v})^{\mathsf{T}}x_{v}$ 3.932M $2.030.7_{\color[rgb]{0.1914,0.3359375,0.9414}{+2.0}}$ $7.299.4_{\color[rgb]{0.1914,0.3359375,0.9414}{+7.2}}$ $2.6645.68_{\color[rgb]{0.1914,0.3359375,0.9414}{+2.66}}$
LoRA $x_{t}(x_{v})^{\mathsf{T}}x_{v}$ 4.746M $1.330.0_{\color[rgb]{0.1914,0.3359375,0.9414}{+1.3}}$ $6.899.0_{\color[rgb]{0.1914,0.3359375,0.9414}{+6.8}}$ $2.3645.38_{\color[rgb]{0.1914,0.3359375,0.9414}{+2.36}}$
@@
$\text{Adapter}^{\dagger}$ $x_{t}(x_{v})^{\mathsf{T}}x_{v}$ 1.830M $0.130.8_{\color[rgb]{0.1914,0.3359375,0.9414}{+0.1}}$ $98.8_{-1.0}$ $45.48_{-0.22}$

### v2

@@ 4.3 Results
$x_{t}\circ x^{\prime}_{v}$ $7.8\mathbf{26.1_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+7.8}}}$ $6.3\mathbf{23.7_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+6.3}}}$ $11.8\mathbf{48.6_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+11.8}}}$ $32.9\mathbf{88.7_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+32.9}}}$ $5.5\mathbf{17.3_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+5.5}}}$ $12.86\mathbf{40.88_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+12.86}}}$ $4.4\underline{25.4_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+4.4}}}$ $2.8\mathbf{23.7_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+2.8}}}$ $4.8\mathbf{48.2_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+4.8}}}$ $17.6\mathbf{87.9_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+17.6}}}$ $3.0\mathbf{17.7_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+3.0}}}$ $6.52\mathbf{40.58_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+6.52}}}$
$x_{t}+x^{\prime}_{v}$ $3.922.2_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+3.9}}$ $2.920.3_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+2.9}}$ $4.241.0_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+4.2}}$ $17.773.5_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+17.7}}$ $3.014.8_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+3.0}}$ $6.3434.36_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+6.34}}$ $4.4\underline{25.4_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+4.4}}}$ $2.2\underline{23.1_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+2.2}}}$ $4.7\underline{47.1_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+4.7}}}$ $16.8\underline{87.1_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+16.8}}}$ $2.317.0_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+2.3}}$ $5.88\underline{39.94_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+5.88}}}$
$x_{t}(x_{v})^{\mathsf{T}}x_{v}$ $6.5\underline{24.8_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+6.5}}}$ $5.2\underline{22.6_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+5.2}}}$ $8.6\underline{45.4_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+8.6}}}$ $\underline{84.9_{+29.1}}$ $5.0\underline{16.8_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+5.0}}}$ $10.88\underline{38.90_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+10.88}}}$ $3.924.9_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+3.9}}$ $1.822.7_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+1.8}}$ $2.045.4_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+2.0}}$ $13.583.8_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+13.5}}$ $2.216.9_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+2.2}}$ $4.6838.74_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+4.68}}$
$x_{t}x^{\prime\prime}_{v}$ $5.623.9_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+5.6}}$ $4.521.9_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+4.5}}$ $7.143.9_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+7.1}}$ $14.780.5_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+14.7}}$ $4.416.2_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+4.4}}$ $9.2637.28_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+9.26}}$ $4.5\mathbf{25.5_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+4.5}}}$ $2.022.9_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+2.0}}$ $2.646.0_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+2.6}}$ $14.985.2_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+14.9}}$ $2.5\underline{17.2_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+2.5}}}$ $5.3039.36_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+5.30}}$
@@
$x_{t}\circ x^{\prime}_{v}$ $8.7\underline{24.6_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+8.7}}}$ $4.6\underline{23.1_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+4.6}}}$ $9.446.4_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+9.4}}$ $22.984.5_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+22.9}}$ $3.1\underline{17.2_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+3.1}}}$ $9.7439.16_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+9.74}}$ $7.523.6_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+7.5}}$ $4.2\underline{22.6_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+4.2}}}$ $8.6\underline{44.8_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+8.6}}}$ $21.3\underline{82.8_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+21.3}}}$ $3.2\underline{17.1_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+3.2}}}$ $8.96\underline{38.18_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+8.96}}}$
$x_{t}+x^{\prime}_{v}$ $5.121.0_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+5.1}}$ $3.021.5_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+3.0}}$ $5.442.4_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+5.4}}$ $14.475.0_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+14.4}}$ $1.916.0_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+1.9}}$ $5.7635.18_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+5.76}}$ $3.219.3_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+3.2}}$ $2.420.8_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+2.4}}$ $4.740.9_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+4.7}}$ $13.875.2_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+13.8}}$ $2.316.2_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+2.3}}$ $5.2634.48_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+5.26}}$
$x_{t}(x_{v})^{\mathsf{T}}x_{v}$ $\underline{26.1_{+10.2}}$ $4.7\mathbf{23.2_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+4.7}}}$ $9.9\mathbf{46.9_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+9.9}}}$ $23.8\underline{85.4_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+23.8}}}$ $3.2\mathbf{17.3_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+3.2}}}$ $10.36\underline{39.78_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+10.36}}}$ $8.0\underline{24.1_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+8.0}}}$ $4.022.4_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+4.0}}$ $8.044.2_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+8.0}}$ $21.282.7_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+21.2}}$ $3.016.9_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+3.0}}$ $8.8438.06_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+8.84}}$
$x_{t}x^{\prime\prime}_{v}$ $10.7\mathbf{26.6_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+10.7}}}$ $4.523.0_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+4.5}}$ $9.8\underline{46.8_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+9.8}}}$ $24.2\mathbf{85.8_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+24.2}}}$ $3.1\underline{17.2_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+3.1}}}$ $10.46\mathbf{39.88_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+10.46}}}$ $9.9\mathbf{26.0_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+9.9}}}$ $5.1\mathbf{23.5_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+5.1}}}$ $10.6\mathbf{46.8_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+10.6}}}$ $24.8\mathbf{86.3_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+24.8}}}$ $3.9\mathbf{17.8_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+3.9}}}$ $10.86\mathbf{40.08_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+10.86}}}$
@@
LoRA $x_{t}(x_{v})^{\mathsf{T}}x_{v}$ 3.932M $2.030.7_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+2.0}}$ $7.299.4_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+7.2}}$ $2.6645.68_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+2.66}}$
LoRA $x_{t}(x_{v})^{\mathsf{T}}x_{v}$ 4.746M $1.330.0_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+1.3}}$ $6.899.0_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+6.8}}$ $2.3645.38_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+2.36}}$
@@
$\text{Adapter}^{\dagger}$ $x_{t}(x_{v})^{\mathsf{T}}x_{v}$ 1.830M $0.130.8_{\color[rgb]{0.1914,0.3359375,0.9414}\definecolor[named]{pgfstrokecolor}{rgb}{0.1914,0.3359375,0.9414}{+0.1}}$ $98.8_{-1.0}$ $45.48_{-0.22}$